OpenStudy (anonymous):

simplify 4x3 - 8x + 16 plz help

OpenStudy (anonymous):

aussy here the real question thats the old one Simplify: 6a(4a - 2) + 3a(a + 7)

mathslover (mathslover):

@cody1221 u r not allowed to insult any1 in os ... u may get suspended

mathslover (mathslover):

that is good spirit dude

OpenStudy (anonymous):

ok you use the distributive property Distribute 6a to everything within the parenthesis 6a(4a-2)= 24a-12a= subtract like terms you should get 12a Then move on to the next part of the equation which is 3a(a+7) and repeat the distributive property which is 3a^2+21a= These are not like terms so they remain the same then you finish your equation 12a+3a^2+21a------.> again you see like terms; 12a+21a=33a >>>> simplified= 3a^2+33a<<<----